You win Rs 10,000 in a horse race. You have a choice between spending the money now and putting it in a 5% FD. What is the opportunity cost of spending the Rs 10,000 now?


1
Expert's answer
2021-10-13T15:45:23-0400

The opportunity cost would be the interest foregone by not investing the money

"Interest= P \times R \times T"

Let use T = 5 years

"= 10000 \times 0.05 \times 5 \\

= 2500"